One year​ ago, your company purchased a machine used in manufacturing for $ 105 comma 000. You have learned that a new machine is available that offers many advantages and you can purchase it for $ 150 comma 000 today. It will be depreciated on a​ straight-line basis over 10 years and has no salvage value. You expect that the new machine will produce a gross margin​ (revenues minus operating expenses other than​ depreciation) of $ 45 comma 000 per year for the next 10 years. The current machine is expected to produce a gross margin of $ 23 comma 000 per year. The current machine is being depreciated on a​ straight-line basis over a useful life of 11​ years, and has no salvage​ value, so depreciation expense for the current machine is $ 9 comma 545 per year. The market value today of the current machine is $ 60 comma 000. Your​ company's tax rate is 38 %​, and the opportunity cost of capital for this type of equipment is 10 %. Should your company replace its​ year-old machine? The NPV of replacing the​ year-old machine is
Answer #1

Answer:

Current Machine:

Book value now = Purchase cost - Depreciation for 1 year = 105000 - $9545 =$95455

Market value now = $60,000

Loss on sale = 95455 - 60000 = $35455

Tax benefit on loss = 35455 * 38% = $13472.90

Sales proceeds including tax benefit = 60000 + 13472.90 = $73,472.90

Year 0 cash out flow:

Initial investment = New machine cost - Sales proceeds of current machine including tax benefit

= 150000 - 73472.90

= $76,527.10

Annual cash flow:

Gross margin new machine = $45,000

Gross margin current machine = $23,000

Incremental margin net of tax =(45000 - 23000) * (1 - 38%) = $13,640

Annual depreciation new machine = 150000 / 10 = $15,000

Incremental depreciation = 15000 - 9545 = $5455

Depreciation tax shield = 5455 * 38% = $2072.90

Annual cash flow = 13640 + 2072.90 = $15,712.90

NPV = Annual cash flow * PV of $1 annuity for 10 years at 10% rate - Initial investment

= 15712.90 * (1 - 1 /(1 +10%) 10) / 10% - 76527.10

= $20021.87

NPV = $20,022

[Note: The depreciation of current machine is taken as $9545 as given in the question. However of we take exact depreciation (105000/ 11=) $9545.45, then NPV will work out at $20,020.64 or $20,021.]